Question 14 (1 point) Suppose that Spain and France both produce ships and grapes, which are sold for the same price in both countries. The table below shows the combinations of the two goods that each country can produce in one year using the same amounts of capital and labor. Suppose that Spain is currently producing 30 ships and 30 tons of grapes, and France is currently producing 30 ships and 30 tons of grapes. If Spain and France each specialize in producing only one good (the good for which each has a comparative advantage), then how many additional tons of grapes can be produced for the two countries combined? Provide your answer as a number rounded to two decimal places. Use to indicate negative amounts. Do not include any symbols, such as S, %, or , in your answer Spain France Shi Shi 0 Grapes (in tons) 75 60 45 30 15 Grapes (in tons) 50 40 30 20 10 10 30 40 50 30 45 60 75
Answer #1

Let us find opportunity cost (OC) as follows.

France can produce maximum 75 ships or maximum 50 grapes.

Spain can produce maximum 50 ships or maximum 75 grapes.

In France,

OC of ship = 50/75 = 0.67 grape

OC of grape = 75/50 = 1.5 ship

In Spain,

OC of ship = 75/50 = 1.5 grape

OC of grape = 50/75 = 0.67 ship

Since Spain can produce grape at a lower OC than France can (0.67 < 1.5), Spain has comparative advantage in grapes and will specialize in grape.

Grape produce after specialization ( by Spain only) = 75 tones

Increase in production of grape = 75 - (30 + 30) = 75 - 60 = 15.00